An archeologist found the remains of an ancient spartans’s shield, which she then placed on a grid. If an arc of the shield passes through A(−6,0), B(2,4) and C(6,0), locate the center of the shield, and the standard equation defining its boundary.
1
Expert's answer
2020-11-08T18:38:10-0500

Since we have 3 points, we will assume that the line's equation is this: "y=ax^2+bx+c" .


We put the coordinates into this equation and we get:

"36a-6b+c=0"

"4a+2b+c=4"

"36a+6b+c=0"


"72a+2c=0"

"c=-36a"

"-6b=0"

"b=0"


"4a-36a=4"

"a=\frac{4}{-32}"

"a=" "-\frac{1}{8}"

"c=-36\cdot(-\frac{1}{8})"

"c=\frac{9}{2}"


So the equation is "y=-\frac{1}{8}(x^2-36)" .


Let's find a highest point of this parabola:

"x=\frac{-b}{2a}=0"

"y(0)=\frac{9}{2}" .


Assuming that this equation that we got represents the upper part of the shield and the lower would go through points (-6,0), (2,-4), (6,0), having its equations as "y=\frac{1}{8}(x^2-36)", we get that the center of the shield would be located right in middle between "(0,\frac{9}{2})" and "(0,-\frac{9}{2})", so it's (0,0).


Answer: "y=-\frac{1}{8}(x^2-36)" - equation (for upper part), (0,0) - center